OpenStudy (anonymous):

The number of solution of log base4 (x-1)=log base2 (x-3) is

OpenStudy (anonymous):

2 values of x

OpenStudy (anonymous):

how 2 answers

OpenStudy (dumbcow):

\[\log_{4}(x-1) = \log_{2}(x-3) \] \[\large x-1 = 4^{\log_{2}(x-3)} \] \[\large x-1 = 2^{2\log_{2}(x-3)}\] \[x-1 = (x-3)^{2}\] now just solve the quadratic...so yes there are 2 solutions

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

thanks

OpenStudy (dumbcow):

yw